### v4

# A Hybrid Systems Model of Feedback Optimization for Linear Systems: Convergence and Robustness

Abstract
Feedback optimization algorithms compute inputs to a system using real-time output measurements, which helps mitigate the effects of disturbances. However, existing work often models both system dynamics and computations in either discrete or continuous time, which may not accurately model some applications. In this work, we model linear system dynamics in continuous time, and we model the computations of inputs in discrete time. Therefore, we present a novel hybrid systems model of feedback optimization. We first establish the well-posedness of this hybrid model and establish completeness of solutions while ruling out Zeno behavior. Then we show the state of the system converges exponentially fast to a ball of known radius about a desired goal state. Next we analytically show that this system is robust to perturbations in (i) the values of measured outputs, (ii) the matrices that model the linear time-invariant system, and (iii) the times at which inputs are applied to the system. Simulation results confirm that this approach successfully mitigates the effects of disturbances.
I Introduction
Many automation tasks require optimizing the behavior of a dynamical system, which often involves solving a planning problem offline. With accurate system models, an optimization problem may be solved in a feedforward configuration to generate a reference that is used to drive the system in question [1, 2, 3, 4, 5, 6, 7]. However, errors in a system model can lead to sub-optimal solutions [1] because the inputs applied to a system may not actually produce the intended outputs.
If models are inaccurate, one alternative approach called “feedback optimization” instead measures system outputs [8, 1, 7, 9, 3] and then uses those measurements to optimize inputs with an in-the-loop optimization algorithm. Feedback optimization has been shown to have several benefits in certain settings: it is robust to inaccurate system models and time-varying parameters, achieves constraint satisfaction with minimal model dependence, and eliminates the need for pre-computed set points or reference signals [1, 3]. This approach has been used, for example, in decentralized settings [10, 11], gradient-based feedback control [9], zeroth-order optimization [1], chemical processes [3], and network congestion control [12].
In existing work, feedback optimization has usually been applied to systems with either (a) continuous-time dynamics and a continuous-time optimization algorithm in the loop or (b) discrete-time dynamics and a discrete-time optimization algorithm in the loop [13, 1, 14]. However, physical systems are often modeled in continuous time and digital computers are naturally modeled in discrete time, which means that practical implementations can produce dynamics that are not captured by (a) or (b).
We seek to show that feedback optimization retains its robustness guarantees with continuous-time dynamics and discrete-time optimization, and we therefore develop a hybrid systems model of it. Related work in [15, 16, 17] considers a continuous-time system and discrete-time computations in a sampled-data setting, though we develop new analytical robustness guarantees by developing a hybrid model in the framework of [18]. We discuss this point more below.
The contributions of this paper are the following:
  • We model feedback optimization as a hybrid system and show that (i) it is free from Zeno behavior and (ii) all maximal solutions are complete (Proposition 1).
  • We bound the distance between the state of the hybrid feedback optimization model and a desired goal state (Theorem 1 and Theorem 2).
  • We analytically prove the robustness of the hybrid model to perturbations by showing that there is bounded difference between the solutions to a perturbed system and the solutions to a nominal system (Theorem 3).
  • We show in simulation that hybrid feedback optimization successfully rejects disturbances (Section VI).
Prior work in [15, 16, 17] combines continuous-time dynamics and discrete-time optimization in a sampled-data feedback optimization setting. The main results in [15, 16] show that the closed-loop system is stable for large enough sample times and that it is practically stable under time-varying disturbances with bounded variation. Results in [17] show global exponential stability of the closed-loop system when inputs to the system change at a fixed rate.
We differ from those prior works by considering a hybrid systems model in the framework of [18], which unlocks robustness analyses that we apply in this work. We use this model to show that hybrid feedback optimization is simultaneously robust to several types of disturbances, including disturbances in the values of measured outputs, errors in the times at which new inputs are applied to the system, and errors in all of the matrices that model the linear time-invariant system that is being controlled. To the best of our knowledge, this paper is the first to analytically prove that a hybrid/sampled-data feedback optimization model is simultaneously robust to all of these disturbances.
The rest of the paper is organized as follows. Section II provides background. Section III gives problem statements and the hybrid feedback optimization model. Section IV derives properties of its solutions. Then Section V establishes the solutions’ convergence and the system’s robustness. Section VI presents simulations, and Section VII concludes.
II Preliminaries
This section gives background on feedback optimization.
II-A Notation
Let $\mathbb{R}$ denote the reals and let $\mathbb{N}$ denote the non-negative integers. For a differentiable function $\Phi:\mathbb{R}^{m}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$, let $\nabla_{u{}}\Phi$ denote the partial derivative with respect to its first argument. The symbol $I_{b}$ denotes the identity matrix of dimension $b$. The $2$-norm of a vector $x$ is denoted $\left\|x\right\|$. For a non-empty, compact, convex set $\mathcal{Z}$, the symbol $\Pi_{\mathcal{Z}}[v]$ denotes the Euclidean projection of a point $v$ onto $\mathcal{Z}$, i.e., $\Pi_{\mathcal{Z}}[v]=\operatorname*{arg\,min}_{z\in\mathcal{Z}}\|v-z\|$. We use $\lambda_{i}(N)$ to denote the $i^{th}$ eigenvalue of a matrix $N$, and we use $\mathfrak{Re}\{a\}$ to denote the real part of a complex number $a$. We also write $\lambda_{min}(N)$ and $\lambda_{max}(N)$ for the smallest and largest (real) eigenvalues of a symmetric matrix $N$, respectively. Given $r\geq 0$, we use $B_{r}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}):=\{x\in\mathbb{R}^{n}:\|x-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\|\leq r\}$ to denote the closed Euclidean ball of radius $r$ about the point $\tilde{x}\in\mathbb{R}^{n}$. For a set $S$ and a point $x$, we use $\|x\|_{S}:=\inf_{s\in S}\|x-s\|$.
II-B Feedback Optimization Background and Setup
Now we review “feedback optimization” as defined in the literature, e.g., [1, 19]. At a high level, this class of problems uses real-time measurements from a dynamical system that are fed into an optimization algorithm in a closed-loop structure. The goal in doing so is to compute inputs that optimize the steady-state behavior of the dynamical system.
To the best of our knowledge there has not been a systematic investigation of the robustness of hybrid feedback optimization for a continuous-time system driven by discrete-time computations. Other existing work has developed hybrid models of optimization algorithms [20, 21, 22], though we differ by developing a hybrid model of feedback optimization. There has also been work on hybrid model predictive control [23], but our aim is different because we implement a hybrid framework for feedback optimization. Work in [15, 16, 17] studies feedback optimization in a sampled-data setting, but our use of a hybrid model lets us derive new analytical robustness guarantees that include robustness to errors in the times at which new inputs are applied, perturbations to the values of sampled outputs, and errors in the plant model.
Suppose we have the linear time invariant (LTI) system
$\begin{split}\dot{x}&=Ax+Bu\\ y&=\Psi x+d,\end{split}$ (1)
where $x\in\mathbb{R}^{n}$ is the system’s state, $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is its input, $\mathcal{U}$ is a non-empty, compact, convex set, and $y\in\mathbb{R}^{p}$ is its output. The vector $d\in\mathbb{R}^{p}$ is a constant, unknown disturbance (e.g., bias), which is a typical component of feedback optimization problem formulations [1, 10, 24, 9, 13]. Such disturbances arise for example in various power systems applications [1, 24, 9, 10, 13].
Assumption 1.
The matrix $A$ is Hurwitz.
Remark 1.
We adopt the “stabilize then optimize” approach [15] in which we suppose that a stabilizing controller has already been applied to the system. If Assumption 1 is not satisfied, then it can be enforced for any stabilizable system by doing pre-feedback with a stabilizing controller.
Assumption 1 ensures that a system will eventually reach steady state, which allows its steady-state behavior to be optimized. From (1) the steady-state input-to-output map is $u\mapsto Hu+d$, where $H:=-\Psi A^{-1}B$.
Remark 2.
While one could envision using an estimator to determine $d$, we are interested in cases in which the matrices $A$, $B$, and $\Psi$ are not known exactly, and an observer may have poor accuracy under these conditions. We also emphasize that knowledge of $H:=-\Psi A^{-1}B$ is used only for analysis and is not required to implement our algorithm.
To optimize the system’s steady-state behavior, one can drive its input and output to a solution of
$\min_{u,y{}}\quad$ $\Phi(u,y{})$ (2a)
subject to $y{}=Hu+d,\,\,u\in\mathcal{U},\,\,y{}\in\mathbb{R}^{p},$ (2b)
where $\mathcal{U}$ is a non-empty, compact, convex set and $\Phi:\mathbb{R}^{m}\times\mathbb{R}^{p}\to\mathbb{R}$ is strongly convex in $(u,y)$. These properties ensure (2a)-(2b) has a unique solution.
By incorporating the constraint $y=Hu+d$ into (2a), one could in principle reduce this problem to
$\min_{u}\quad$ $\tilde{\Phi}(u):=\Phi(u,Hu+d)$ (3)
subject to $u\in\mathcal{U}.$ (4)
However, the problem in (3) cannot be solved in practice because the substitution $y=Hu+d$ would require exact knowledge of the disturbance $d$, which may not be available. Instead, feedback optimization is used to repeatedly measure $y$ (which is perturbed by $d$) and then optimize over $u$. In this work, $y$ is sampled at discrete instants of time, and we use $y_{s}$ to denote its sampled value.
Remark 3.
The underlying LTI system need not always be at steady state, but we will use a standard technique in the feedback optimization literature to approximate its outputs as coming from a system at steady state [8, 9, 10]. Mathematically, we approximate a sampled output $y_{s}$ as coming from the steady-state map $y_{s}=Hu+d$ when the input to the system is $u$. This approximation is justified, for example, when the dynamics of the system converge sufficiently quickly. Our results in Sections IV and V still analyze the dynamics for the system, and they use this approximation only to relate outputs to inputs.
We study a closed-loop system that connects the LTI system with a gradient descent algorithm that computes the input $u$. With a sampled output $y_{s}$, the optimization update law is $u_{k+1}=\Pi_{\mathcal{U}}\big[u_{k}-\gamma\nabla_{u}\Phi(u_{k},y_{s})\big]$, where $u_{k}$ is the $k^{th}$ iterate of the gradient descent algorithm.
Then the closed-loop interconnected LTI system is
Plant: $\begin{cases}\dot{x}&=Ax+Bu\\ y&=\Psi x+d,\end{cases}$ (5)
Controller: $\begin{cases}u_{k+1}=\Pi_{\mathcal{U}}\big[u_{k}-\gamma\nabla_{u}\Phi(u_{k},y_{s})\big].\end{cases}$ (6)
We will formulate and analyze a hybrid model for this interconnection. Neither the changes in the input nor the measurements of the output will be assumed to occur periodically, but instead both can occur sporadically.
II-C Background on Hybrid Systems
In this paper, a hybrid system $\mathcal{H}$ takes the form
$\mathcal{H}=\begin{cases}\begin{aligned} \dot{\zeta}&\in F(\zeta)\quad&\zeta\in C\\ \zeta^{+}&\in G(\zeta)\quad&\zeta\in D\end{aligned}\end{cases},$ (7)
where $\zeta\in\mathbb{R}^{n}$ is the system’s state vector and the maps $F$ and $G$ are set valued in general. The function $F$ defines the flow map and governs the continuous dynamics within the flow set $C$, while $G$ defines the jump map, which models the system’s discrete behavior within the jump set $D$.
Definition 1 (Hybrid Basic Conditions [18]).
A hybrid system $\mathcal{H}$ with data $(C,F,D,G)$ satisfies the hybrid basic conditions if
  1. 1.
    $C$ and $D$ are closed subsets of $\mathbb{R}^{n}$;
  2. 2.
    $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is outer semicontinuous [^1] A set-valued mapping $M:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{n}$ is outer semicontinuous (osc) at $x\in\mathbb{R}^{m}$ if for every sequence of points $\{x_{i}\}_{i\in\mathbb{N}}$ convergent to $x$ and any convergent sequence of points $\{y_{i}\}_{i\in\mathbb{N}}$ with $y_{i}\in M(x_{i})$, one has $y\in M(x)$, where $\lim_{i\rightarrow\infty}y_{i}=y$ [18]., and locally bounded [^2] A set-valued mapping $M:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{n}$ is locally bounded at $x\in\mathbb{R}^{m}$ if there is a neighborhood $U_{x}$ of $x$ such that $M(U_{x})\subset\mathbb{R}^{n}$ is bounded [18]. relative to $C$, $C\subset\text{dom}\penalty 10000\ F,$ and $F(\zeta)$ is convex for every $\zeta\in C$
  3. 3.
    $G:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is outer semicontinuous and locally bounded relative to $D$, and $D\subset\text{dom}\penalty 10000\ G$.
If a hybrid system satisfies the hybrid basic conditions, then it is well-posed by [18, Theorem 6.30]. We use this property in Section V to show that errors in the models of $F$ and $G$ up to a certain threshold produce bounded changes in the resulting closed-loop system trajectories over compact time horizons. It is not automatic to formulate a well-posed hybrid system, and this paper will do so for a hybrid model of feedback optimization.
For a hybrid system $\mathcal{H}$, its solutions, denoted by $\phi$, are hybrid arcs that can in general be maximal [^3] A solution $\phi$ to $\mathcal{H}$ is maximal if there does not exist another solution $\psi$ to $\mathcal{H}$ such that dom $\phi$ is a proper subset of dom $\psi$ and $\phi(t,j)=\psi(t,j)$ for all $(t,j)\in\textnormal{dom }\phi$ [18]., complete [^4] The solution $\phi$ is complete if $\textnormal{dom }\phi$ is unbounded, i.e., if $\textnormal{length}(\textnormal{dom }\phi)=\sup_{t}\textnormal{dom }\phi+\sup_{j}\textnormal{dom }\phi=\infty$ [18]., and Zeno [^5] The solution $\phi$ is Zeno if it is complete and $\sup_{t}\textnormal{dom }\phi<\infty$ [18].. Complete solutions are defined over arbitrarily long time horizons, and Zeno behavior implies that solutions undergo an infinite number of jumps in finite time. Zeno behavior implies that a system’s states stop flowing in finite time, which we will rule out for feedback optimization.
III A Hybrid Model for Feedback Optimization
The problems we solve in this paper are as follows.
Problem 1.
Formulate a well-posed hybrid feedback optimization model of the plant and controller in (5) that captures (i) how outputs are intermittently measured and used in the optimization-based controller, and (ii) how inputs are computed and applied to the system.
Problem 2.
Prove the convergence of solutions toward a desired goal state and bound their steady-state error.
Problem 3.
Show that the hybrid feedback optimization model is robust to perturbations in the LTI system model, measurements of outputs, and the times at which inputs are applied to the LTI system, in the sense that these perturbations induce bounded changes in solutions.
III-A Overview of Hybrid Feedback Optimization
The continuous-time system in (5) receives inputs from the discrete-time optimization algorithm in (5), and those inputs only change at certain instants of time. Between these changes, inputs applied to the system are held constant. Similarly, the optimization algorithm measures an output of the system and uses it to perform some number of computations to optimize inputs. This sampled value of the output is held constant while optimizing an input, and it does not change until a new output is sampled.
The flow map $F$ from Definition 1 will model the LTI system dynamics in (1) with piecewise constant inputs. The jump map $G$ from Definition 1 will model both the sampling of outputs and the application of new inputs to the system. It will also model computations done by the optimization algorithm, which are generated at discrete points in time between the times at which the output value is sampled.
III-B Optimization Problem Setup
We consider objectives $\Phi$ of the form
$\Phi(u,y{})=\frac{1}{2}u^{\top}Q_{u{}}u+\frac{1}{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(y_{s}-\hat{y})}^{\top}Q_{y{}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(y_{s}-\hat{y})}$, (8)
where $Q_{u{}}\in\mathbb{R}^{m\times m}$ and $Q_{y{}}\in\mathbb{R}^{p\times p}$ are symmetric and positive definite, and $\hat{y}\in\mathbb{R}^{p}$ is a constant, desired output value that is user-specified.
Quadratic objectives have been widely used in the feedback optimization literature [1, 25, 10, 13, 11, 19], and we emphasize that our hybrid model is not restricted to using quadratic objectives. Instead, in this initial work we focus on quadratic objectives to develop a hybrid systems model for feedback optimization, and we defer the study of other objectives (including nonconvex objectives) to future work that will build on the present paper.
III-C Hybrid Modeling and Flow and Jump Sets
The state of the hybrid system includes $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{m}$ which are, respectively, the state and input of the LTI system in (1). It also includes the vector $y_{s}\in\mathbb{R}^{p}$, which is the value of the sampled output of the LTI system that is used in the underlying optimization algorithm, $z\in\mathbb{R}^{m}$, which is the current iterate of that optimization algorithm, $\tau_{c}\in\mathbb{R}$, which is a timer that tracks the amount of continuous time left until the input to the system changes, and $\tau_{g}\in\mathbb{R}$, which is a timer that accounts for the amount of continuous time needed to complete an iteration of the optimization algorithm. The full state is
$\zeta:=\left(\begin{array}[]{cccccc}\!x^{\top}&u^{\top}&y_{s}^{\top}&z^{\top}&\tau_{c}&\tau_{g}\end{array}\!\right)^{\top}\in\mathcal{X}:=\mathbb{R}^{n+2m+p+2}.$ (9)
The timers $\tau_{c}$ and $\tau_{g}$ count down from some positive numbers to zero, and jumps occur only when they reach zero. The state is allowed to flow while both $\tau_{c}>0$ and $\tau_{g}>0$, and the state stops flowing and undergoes a jump when $\tau_{c}=0$ and/or $\tau_{g}=0$. These conditions are captured by the flow set $C$ and jump set $D$, defined as
$C$ $:=\big\{\zeta\in\mathcal{X}\mid\tau_{c}\in[0,\tau_{c,\max}],\tau_{g}\in[0,\tau_{g,comp}]\big\}$ (10)
$D$ $:=\big\{\zeta\in\mathcal{X}\mid\tau_{c}=0\textnormal{ or }\tau_{g}=0\big\},$ (11)
where $\tau_{c,max}>0$ is the maximum amount of time between changes in the input and $\tau_{g,comp}>0$ is the amount of time needed to perform a gradient descent iteration.
III-D Flow Map Definition
The flow map is derived from (1), which defines $\dot{x}$. We note that $y_{s}{}$, the sampled output used by the optimization algorithm, does not vary continuously because it is measured at certain time instants and is held constant between measurements. The timers $\tau_{c}$ and $\tau_{g}$ count down to zero continuously and with unit rate, while all other states only change during jumps. Therefore, the flow map is
$\dot{\zeta}=F(\zeta):=\left(\begin{array}[]{c}Ax+Bu\\ 0\\ 0\\ 0\\ -\hskip-2.0pt1\\ -\hskip-2.0pt1\end{array}\right).$ (12)
III-E Jump Map Definition
The jump map has three cases: (i) $\tau_{g}=0$ with $\tau_{c}>0$, (ii) $\tau_{c}=0$ with $\tau_{g}>0$, and (iii) $\tau_{c}=\tau_{g}=0$.
In case (i), a single gradient descent step has been completed, but since $\tau_{c}>0$, the iterate $z$ is not applied as the system input. The jump map for this case updates the state $z$ using a gradient descent step of the form $z^{+}=\Pi_{\mathcal{U}}\big[z-\gamma\nabla_{u}\Phi(z,y_{s})\big]$, where $\gamma>0$ is a stepsize. Here $y_{s}\in\mathbb{R}^{p}$ is the most recently sampled value of the system output. The jump map resets $\tau_{g}$ to $\tau_{g,comp}$ so that the computation of a new gradient descent iteration can begin. Other states jump to their current values, which leaves them unchanged. The jump map for this case is
$G_{1}(\zeta)=\left(\begin{array}[]{c}x\\ u\\ y_{s}\\ \Pi_{\mathcal{U}}\big[z-\gamma\nabla_{u}\Phi(z,y_{s})\big]\\ \tau_{c}\\ \tau_{g,comp}\end{array}\right).$ (13)
In case (ii), when $\tau_{c}=0$ and $\tau_{g}>0$, a new input is applied to the system, a new output is sampled, and the timer $\tau_{c}$ resets to some point in the interval $\tau_{c,min},\tau_{c,max}]$, where $0<\tau_{c,min}\leq\tau_{c,max}$. This range of times represents indeterminacy in the amount of time that elapses between the application of successive inputs to the system.
Assumption 2.
There exists $\ell\in\mathbb{N}$ with $\ell\geq 1$ such that $\ell\tau_{g,comp}\leq\tau_{c,\min}$.
Assumption 2 ensures that, when the system is properly initialized, at least $\ell$ gradient descent iterations are performed between any consecutive changes in the input. This condition is a mild form of timescale separation and it will be used later in Theorem 1. When the input changes, it is set equal to $z$. When $\tau_{c}$ reaches 0, the LTI system output is sampled and stored in $y_{s}$, which is held constant until the next sample. The states $x$ and $z$ do not change, and the jump map is
$G_{2}(\zeta)=\left(\begin{array}[]{c}x\\ z\\ Hu+d\\ z\\ {[\tau_{c,min},\tau_{c,max}]}\\ \tau_{g}\end{array}\right),$ (14)
where, as described in Remark 3, we approximate the output $y_{s}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Psi}x+d$ with $y_{s}=Hu+d$.
In case (iii), where both $\tau_{c}=0$ and $\tau_{g}=0$, we combine cases (i) and (ii), and the system executes $G_{1}$ and then $G_{2}$ or $G_{2}$ and then $G_{1}$. The full jump map $G$ is defined as
[…]
where $G_{3}(\zeta)=G_{1}(\zeta)\cup G_{2}(\zeta)$. The jump set $D$ is equal to
$D=D_{1}\cup D_{2},$ (16)
where $D_{1}=\{\zeta\in\mathcal{X}:\tau_{g}=0\}$ and $D_{2}=\{\zeta\in\mathcal{X}:\tau_{c}=0\}$.
The full hybrid model of feedback optimization is
$\mathcal{H}_{FO}:=(C,F,D,G),$ (17)
where $C$ is from (10), $F$ is from (12), $D$ is from (16), and $G$ is from (15).
IV Properties of Hybrid Feedback Optimization
In this section, we show $\mathcal{H}_{FO}$ satisfies certain technical conditions that ensure its solutions exist for all time, which completes our solution to Problem 1.
IV-A Well-Posedness and Existence of Solutions
Toward establishing that solutions to $\mathcal{H}_{FO}$ are defined for all time, we have the following.
Lemma 1.
The hybrid feedback optimization model $\mathcal{H}_{FO}$ in (17) is well-posed in the sense that it satisfies Definition 1.
Proof.
By inspection, the set $C$ in (10) and the set $D$ in (16) together satisfy Condition 1 in Definition 1. The map $F$ in (12) is defined everywhere on $C$ and outputs a singleton that is a linear function of the state, and thus it satisfies Condition 2 in Definition 1.
To show that the map $G$ in (15) satisfies Condition 3, we can use Lemma 4 in Appendix -A. The jump map $G_{1}$ in (13) is outer semicontinuous because the projection mapping $\Pi_{\mathcal{U}}[\cdot]$ is continuous and because $G_{1}$ outputs a singleton. The jump map $G_{2}$ in (14) is outer semicontinuous because its only set-valued entry outputs a compact interval and all other entries are singletons. Then the feedback optimization jump map in (15) has the structure of the jump map in Lemma 4, and the feedback optimization jump set in (16) has the same structure as the jump set in Lemma 4. Therefore, using Lemma 4, we see that the feedback optimization jump map $G$ in (15) is both outer semicontinuous and locally bounded relative to the closed set $D$. Then Condition 3 of Definition 1 is satisfied. Therefore, all conditions of Definition 1 are satisfied, and the system $\mathcal{H}_{FO}$ is well-posed. ∎
The following result shows that all maximal solutions to the system $\mathcal{H}_{FO}$ are complete and non-Zeno.
Proposition 1 (Completeness of Maximal Solutions).
Consider the hybrid feedback optimization model $\mathcal{H}_{FO}$ from (17). From every point in $C\cup D$ there exists a nontrivial solution. All maximal solutions are complete and non-Zeno.
Proof.
Since $\mathcal{H}_{FO}$ satisfies Definition 1, we can apply Lemma 5 in Appendix -B to establish the claim. First we will show that condition (VC) holds for all $\D\nu\in C\backslash$ D. Consider an arbitrary $\D\nu\in C\backslash$ D, and let $U$ be a neighborhood of $\nu$. We wish to show that $F(\zeta)\cap T_{C}(\zeta)\neq\emptyset$ for $\zeta\in U\cap C$, where $T_{C}(\zeta)$ is the tangent cone of the set $C$ at the point $\zeta$. Using the flow map $F$ from (12), we see that only $\dot{x}$, $\dot{\tau_{c}}$, and $\dot{\tau_{g}}$ are non-zero during flows. In addition, $C$ does not restrict $x$, which implies that $x$ can flow in any direction at any time and remain feasible. Conversely, the timers $\tau_{c}$ and $\tau_{g}$ take values in compact intervals, which implies that some directions are infeasible at some points in time. Therefore, the satisfaction of the condition $F(\zeta)\cap T_{C}(\zeta)\neq\emptyset$ is determined by the dynamics of $\tau_{c}$ and $\tau_{g}$.
To show that $F(\zeta)\cap T_{C}(\zeta)\neq\emptyset$, we compute the tangent cone as
$T_{C}($ $\zeta):=\penalty 10000\$(18)
[…]
By inspection, it holds that $F(\zeta)\in T_{C}(\zeta)$ for every $\D\zeta\in C\backslash$ D and condition (VC) from Lemma 5 is satisfied for every $\D\nu\in C\backslash$ D. Then there exists a nontrivial solution $\phi$ to $\mathcal{H}_{FO}$ with $\phi(0,0)=\nu$. Let $\mathcal{S}_{\mathcal{H}_{FO}}$ denote the set of all such solutions. Every solution $\phi\in\mathcal{S}_{\mathcal{H}_{FO}}$ satisfies one of the three conditions of Lemma 5.
By inspection, we have $G(D)\subset C\cup D$, which implies that $3)$ in Lemma 5 does not occur. Regarding $2)$, during flows, the components of the solution that change are $x,\tau_{c}$, and $\tau_{g}$, and thus only their boundedness needs to be verified. Since $\tau_{c}$ and $\tau_{g}$ take values in compact sets, they are bounded and cannot blow up to infinity. And since $u\in\mathcal{U}$, which is a compact set, we see that $u$ is bounded, and there exists some finite $u_{max}\geq 0$ such that $\|u\|\leq u_{max}$. Since the mapping $(x,u)\mapsto Ax+Bu$ is Lipschitz, $F$ in (12) is globally Lipschitz. This property and the boundedness of inputs together imply that Condition $2)$ from Lemma 5 does not hold. Therefore, Condition $1)$ from Lemma 5 does hold, and all maximal solutions to $\mathcal{H}_{FO}$ are complete.
Resets in $G$ are triggered only when either one of the timers has reached zero, and $G$ resets any timers that have reached zero to non-zero values. Then $G(D)\cap D=\emptyset$, which rules out Zeno behavior by Proposition 2.34 in [26]. Then all maximal solutions are complete and non-Zeno. ∎
IV-B Algorithm Framework
Consider an initial condition $\nu$ that satisfies
$\tau_{c}(0,0)\in[\tau_{c\min},\tau_{c,\max}]\textnormal{ and }\tau_{g}(0,0)=\tau_{g,comp},$ (20)
and consider a solution $\phi$ to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)=\nu$. We use $\alpha(i)$ to denote the number of gradient descent iterations that are performed when computing the $(i+1)^{th}$ value of the input $u$. Assumption 2 implies that $\alpha(i)\geq\ell\geq 1$ for all $i\in\mathbb{N}$.
After the initial input to the system $u(0,0)$ is applied, the system performs $\alpha(0)$ gradient descent iterations (which are $\alpha(0)$ Case (i) jumps) before the input to the LTI system is changed. The initial condition for these iterates is $z_{0}(0,0)=u(0,0)$, and the $k^{th}$ such iterate is denoted $_{k}(t_{k},k); zz$ without a subscript always denotes a state of $\mathcal{H}_{FO}$, and subscripts are used only in our analyses to keep track of how many iterations have been completed when computing the next input to the LTI system. Before $u$ jumps for the first time, $\alpha(0)$ iterates are computed, the last of which is denoted $z_{\alpha(0)}\big(t_{\alpha(0)},\alpha(0)\big)$.
Figure 1: A visual representation of the input evolution of $\mathcal{H}_{FO}$. There are $\alpha(p)$ iterations of gradient descent when computing the $(p+1)^{\textnormal{th}}$ input. The input is changed at hybrid times of the form $(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p)$.
[…]
We can identify the general pattern that occurs at an arbitrary case (i) jump. Suppose that $p$ total Case (ii) jumps have occurred so far, which means that the input has changed $p$ times. Then, when a Case (i) jump occurs, it computes an optimization iterate of the form
$z_{k+1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p+k+1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p+k+1)=\\ \Pi_{\mathcal{U}}\big[z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p+k},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p+k)\\ -\gamma\nabla_{u}\Phi\big(z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p+k},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p+k)\big)\big],$ (21)
where for notational simplicity we have used
[…]
For a Case (ii) jump, the $(p+1)^{th}$ input to the system is
$u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p)=\\ z_{\alpha(p-1)}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+\alpha(p-1)+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+\alpha(p-1)+p-1),$ (23)
which indicates that the new value of the input is equal to the most recently computed optimization iterate. The initial iterate for computing the next input is $z_{0}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p)$, and it is set equal to the same value as $u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p)$.
V Convergence Analysis
[…]
V-A States with Piecewise Constant Inputs
The next result computes the difference between the state $x$ and its optimal steady-state value $\tilde{x}$.
Lemma 2.
[…]
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\\ \cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}$, (24)
where we define $t_{0}:=0$, the integer $P=\max\{p\in\mathbb{N}:\bar{\alpha}(p)+p\leq j\}$ is the number of times the input to the LTI system has changed before hybrid time $(t,j)$, and $x(0,0)$ is a component of the initial condition $\phi(0,0)$.
Proof.
The model of $\mathcal{H}_{FO}$ applies piecewise constant inputs to the underlying LTI system. We will therefore integrate the underlying LTI dynamics over the intervals across which the input is constant and compute the difference between $x$ and $\tilde{x}$. For $p\in\mathbb{N}$, the input is equal to $u(t_{\bar{\alpha}(p)+p},\bar{\alpha}(p)+p)$ over intervals of the form $t_{\bar{\alpha}(p)+p},t_{\bar{\alpha}(p+1)+p+1}]$ and therefore we may write
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})+\int_{0}^{t}e^{A(t-\tau)}Bu(\tau)d\tau-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\\ =e^{At}x({0},{0})+\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\\ \cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot Bu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(P\right)}+P})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}$. (25)
V-B Input Convergence
The following lemma relates successive iterates that are used to the compute the inputs to the LTI system. In it, we use
$L=\lambda_{max}(Q_{u}+H^{\top}Q_{y}H).$ (26)
Lemma 3 (Input Convergence Rate).
Consider the hybrid system $\mathcal{H}_{FO}$ from (17) and consider an objective of the form of (8). Suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$. Let $\phi$ denote a maximal solution to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)=\nu$ that satisfies (20). For any $(t,j)\in\textnormal{dom }\phi$, set $P=\max\{p\in\mathbb{N}:\bar{\alpha}(p)+p\leq j\}$. Then, for any integer $p\in\{0,\ldots,P\}$ the state $z$ obeys
$\left\|z_{\alpha({p})}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+\alpha({p})+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+\alpha({p})+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|\\ \leq q^{\frac{\alpha({p})-1}{2}}\left\|z_{1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|,$ (27)
where
$z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})=\operatorname*{arg\,min}_{u\in\mathcal{U}}\Phi\big(u,y_{s}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p)\big)$ (28)
and $q:=1-2\gamma\lambda_{min}(Q_{u})+\gamma^{2}L^{2}\in\left(0,1\right),$ where $L$ is from (26).
Proof.
The steps of the proof follow those of a standard proof in the convex optimization literature for the minimization of a strongly convex function, though we present the proof in the hybrid context. For the hybrid system $\mathcal{H}_{FO}$, we can quantify convergence of the computation of inputs by examining the distance of an intermediate iterate from its optimal value after $k+1$ iterations of gradient descent. That is, we can bound the term
$\left\|z_{k+1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2},$ (29)
where $k+1$ is some number of iterations between $1$ and $\alpha({p})$. We can then express $z_{k+1}$ in terms of $z_{k}$ and use the fact that
$z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})=\Pi_{\mathcal{U}}\Big[z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\\ -\gamma\nabla_{u}\Phi\big(z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}),y_{s}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p)\big)\Big],$ (30)
i.e., $z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})$ is a fixed point of the projected gradient descent update law. Doing so gives
$\left\|z_{k+1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ =\left\|\Pi_{\mathcal{U}}\big[z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-\gamma\nabla_{u}\Phi\left(z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right)\big]\right.\\ -\Pi_{\mathcal{U}}\big[z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\\ \left.-\gamma\nabla_{u}\Phi\left(z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right)\big]\right\|^{2},$ (31)
where for ease of notation we have used
$\nabla_{u}\Phi\left(z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right):=\\ \nabla_{u}\Phi(z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p}),y_{s}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})),$
and similar for $\nabla_{u}\Phi\left(z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right)$. The non-expansive property of $\Pi_{\mathcal{U}}$ lets us remove the projections and attain an upper bound. Doing this and expanding gives
$\left\|z_{k+1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ \leq\left\|z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ -2\gamma\left(z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right)^{\top}\cdot\\ \left(\nabla_{u}\Phi\left(z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right)\right.\\ \left.-\nabla_{u}\Phi\left(z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right)\right)\\ +\gamma^{2}\left\|\nabla_{u}\Phi\left(z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right)\right.\\ \left.-\nabla_{u}\Phi\left(z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right)\right\|^{2}.$ (32)
Using the $L$-Lipschitz property of $\nabla\Phi_{u}$ from (26), we find
$\left\|z_{k+1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ \leq\left\|z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ -2\gamma\left(z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right)^{\top}\cdot\\ \left(\nabla_{u}\Phi\left(z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right)\right.\\ \left.-\nabla_{u}\Phi\left(z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right)\right)\\ +\gamma^{2}L^{2}\left\|z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}.$ (33)
Then, since $Q_{u}$ is symmetric and positive definite, the function $\Phi(\cdot,y_{s}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}))$ is $\lambda_{min}(Q_{u})$-strongly convex, and its gradient is $\lambda_{min}(Q_{u})$-strongly monotone. Then we find
$\left\|z_{k+1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ \leq\left\|z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ -2\gamma\lambda_{min}(Q_{u})\left\|z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ +\gamma^{2}L^{2}\left\|z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2},$ (34)
which simplifies to
$\left\|z_{k+1}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+1+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ \leq\left(1-2\gamma\lambda_{min}(Q_{u})+\gamma^{2}L^{2}\right)\left\|z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}\\ =q\left\|z_{k}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+k+p})\right.\\ \left.-z^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|^{2}.$ (35)
To ensure that $q\in\left(0,1\right)$, we use [27, Theorem 2.1.15], which shows that $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$ gives $q\in(0,1)$. Then iteratively applying (35) and taking the square root completes the proof. ∎
In Lemma 3 the term $q^{\frac{\alpha({p})-1}{2}}$ shows why the condition $\ell\tau_{g,comp}\leq\tau_{c,min}$ is required in Assumption 2. If we did not enforce that condition, then the exponent of $q$ could be negative, in which case the optimization algorithm could diverge.
V-C Complete Hybrid Convergence
The next result is our first main result. It bounds the distance from the state of the underlying LTI system, namely $x$, to its optimal steady-state value, $\tilde{x}$. To state this result, we define $\mathcal{Y}=\{y\in\mathbb{R}^{p}:y=Hu+d,u\in\mathcal{U}\}$, which is compact due to $\mathcal{U}$ being compact. Mathematically, we bound the distance between a solution $\phi$ of $\mathcal{H}_{FO}$ and the set
$\mathcal{A}:={B_{r}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})}\times\mathcal{U}\times\mathcal{Y}\times\mathcal{U}\times[0,\tau_{c,\max}]\times[0,\tau_{g,comp}],$ (36)
where
$r=\frac{M\left\|B\right\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{\mathcal{U}}}}{\rho}(1+q^{\frac{\ell}{2}}),$ (37)
$M\geq 1$ is a constant,
$\rho=\min_{i\in\{1,\ldots,n\}}|\mathfrak{Re}\{\lambda_{i}(A)\}|,$ (38)
[…]
Theorem 1 (Complete Hybrid Convergence).
Consider the hybrid system $\mathcal{H}_{FO}$ from (17) and suppose that Assumptions 1 and 2 hold. Consider objectives of the form of (8), and suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$, where $Q_{u}$ is from (8) and $L:=\lambda_{max}(Q_{u}+H^{\top}Q_{y}H)$. For each maximal solution $\phi$ to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)=\nu$ that satisfies (20), for each $(t,j)\in\textnormal{dom }\phi$,
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)\|\phi(0,0)\|_{\mathcal{A}}\\ +(M-1)\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big(1+q^{\frac{\ell}{2}}\big)\exp({-\rho}t),$ (39)
where $d_{\mathcal{U}}=\max_{u_{1},u_{2}\in\mathcal{U}}\|u_{1}-u_{2}\|$ is the diameter of the set $\mathcal{U}$, $\rho$ is from (38), $\ell\geq 1$ is from Assumption 2, $\mathcal{A}$ is from (36), $q:=1-2\gamma\lambda_{min}(Q_{u})+\gamma^{2}L^{2}\in\left(0,1\right)$, and $M\geq 1$ is a constant. In particular, each such solution satisfies
$\lim_{t+j\rightarrow\infty}\left\|\phi(t,j)\right\|_{\mathcal{A}}=0.$ (40)
Proof.
By definition of $\mathcal{A}$, only the state $x$ in $\phi$ affects the value of $\|\phi(t,j)\|_{\mathcal{A}}$. That is, we have
$\|\phi(t,j)\|_{\mathcal{A}}=\|x(t,j)\|_{B_{r}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}})}$, (41)
and we therefore focus our analysis on $x$. We define $P=\max\{p\in\mathbb{N}:\bar{\alpha}(p)+p\leq j\}$. Using (23), we observe that for each $p\in\{0,\ldots,P\}$ we have
$u{}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\\ =z_{\alpha({p-1})}\big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+\alpha({p-1})+p-1},\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+\alpha({p-1})+p-1\big).$ (42)
The iterates that are computed to give that input are working towards the optimizer
$u^{*}{}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})=z^{*}\Big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\Big),$ (43)
which is defined as
$u^{*}{}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\\ =\operatorname*{arg\,min}_{u\in\mathcal{U}}\Phi\Big(u,y_{s}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1})\Big).$ (44)
We then find that
[…]
where we have taken the following steps. First, we have applied Lemma 3, and then we have expanded the gradient descent law as
$z_{1}\big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p\big)\\ =\Pi_{\mathcal{U}}\bigg[z_{0}\big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\big)\\ -\gamma\nabla_{u}\Phi\Big(z_{0}\big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\big),\\ y_{s}\big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\big)\Big)\Big)\bigg].$ (46)
Next, we have used the fact that the optimum is a fixed point of projected gradient descent, namely
$z^{*}\big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\big)\\ =\Pi_{\mathcal{U}}\Big[z^{*}\Big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\big)\\ -\gamma\nabla_{u}\Phi\Big(z^{*}\big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\big),\\ y_{s}\big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\big)\Big)\Big],$ (47)
and then applied the non-expansive property of $\Pi_{\mathcal{U}}$ to attain an upper bound by removing $\Pi_{\mathcal{U}}$. Then we have approximated the sampled output as described in Remark 3 via
$y_{s}\big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\big)\Big)\\ =Hu(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1)+d\\ =Hz_{0}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1)+d.$ (48)
Next, combining like terms in (45) and applying the triangle inequality gives
$\left\|u{}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|\\ \leq q^{\frac{\alpha({p-1})-1}{2}}\left\|z_{0}\Big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\Big)\right.\\ \left.-z^{*}\Big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\Big)\right\|\\ \cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|.$ (49)
Using
$\left\|z_{0}\Big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\Big)\right.\\ \left.-z^{*}\Big(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p-1\right)}+p-1\Big)\right\|\leq d_{\mathcal{U}}$ (50)
then gives the bound
$\left\|u{}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|\\ \leq q^{\frac{\alpha({p-1})-1}{2}}d_{\mathcal{U}}\cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|.$ (51)
Next, using Lemma 2 and adding and subtracting $u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)$ we have
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot\\ B\big(u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\big)\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau Bu^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}$. (52)
Using $\tilde{x}=-A^{-1}B\tilde{u}$ and adding $\tilde{u}+\tilde{u}$ gives
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot\\ B\big(u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\big)\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-\tilde{u}\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau B(u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{u})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau B\tilde{u}\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau B\tilde{u}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}$. (53)
We know that
$\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau=\int_{0}^{t}e^{A(t-\tau)}d\tau$ (54)
and that
$\int_{0}^{t}e^{A(t-\tau)}d\tau=(e^{At}-I)A^{-1}.$ (55)
These relations allow for simplifying (53) to
$x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}=e^{At}x({0},{0})\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau\cdot\\ B\big(u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\big)\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}d\tau\cdot\\ B\big(u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-\tilde{u}\big)\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}e^{A(t-\tau)}d\tau B(u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{u})\\ +e^{At}A^{-1}B\tilde{u}$. (56)
[…]
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|\left\|u(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})\right\|\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|\left\|u(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)\right\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|\left\|u{}^{*}(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p})-\tilde{u}\right\|\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|\left\|u^{*}(t_{\bar{\alpha}(P)+P},\bar{\alpha}(P)+P)-\tilde{u}\right\|$. (57)
Using (51) and the definition of $d_{\mathcal{U}}$ gives
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|q^{\frac{\alpha({p-1})-1}{2}}d_{\mathcal{U}}\cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\cdot\\ \left\|B\right\|q^{\frac{\alpha({p-1})-1}{2}}d_{\mathcal{U}}\cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\sum^{P-1}_{p=0}\int_{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p\right)}+p}}^{t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}}\left\|e^{A(t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\left(p+1\right)}+p+1}-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}\\ +\int_{t_{\bar{\alpha}(P)+P}}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}.$ (58)
Assumption 2 ensures that $\alpha(p)\geq\ell$ for all $p$. Using this fact and $q\in(0,1)$, we have
$q^{\frac{\alpha(p-1)-1}{2}}\leq q^{\frac{\ell-1}{2}}.$ (59)
Then using (59) in (58) and combining sums of integrals into one integral as was done below (53), we have
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|=\left\|e^{At}\right\|\left\|x({0},{0})-\tilde{x}\right\|\\ +\int_{0}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|q^{\frac{\ell-1}{2}}d_{\mathcal{U}}\\ \cdot\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\\ +\int_{0}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\left\|B\right\|d_{\mathcal{U}}.$ (60)
Next, we observe that
$\int_{0}^{t}\left\|e^{A(t-\tau)}\right\|d\tau\leq M\int_{0}^{t}\exp\big({-\rho}(t-\tau)\big)d\tau\\ =\frac{M}{\rho}\big[1-\exp\big({-\rho}t\big)\big],$ (61)
where the inequality follows from Theorem 2 in [28, Chapter 1.9].
Substituting (61) into (60), using
$\|e^{At}\|\leq M\exp({-\rho}t),$ (62)
and rearranging terms gives
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|\leq M\exp({-\rho}t)\left\|x({0},{0})-\tilde{x}\right\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|q^{\frac{\ell-1}{2}}d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\cdot\\ \|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|.$ (63)
We find that
$\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|\leq q^{\frac{1}{2}},$ (64)
which follows from observing that
$\|I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)\|^{2}\\ =\sup_{x:\|x\|=1}\|[I_{m}-\gamma(Q_{u}+H^{T}Q_{y}H)]x\|^{2}\\ \leq 1-\inf_{x:\|x\|=1}2\gamma x^{T}(Q_{u}+H^{T}Q_{y}H)x\\ +\gamma^{2}\|Q_{u}+H^{T}Q_{y}H\|^{2}.$ (65)
Then, since $H^{T}Q_{y}H\succeq 0$, we have
$\inf_{x:\|x\|=1}2\gamma x^{T}(Q_{u}+H^{T}Q_{y}H)x\\ \geq\inf_{x:\|x\|=1}2\gamma x^{T}Q_{u}x=2\gamma\lambda_{min}(Q_{u}).$ (66)
We use (66) and (26) in (65), and then taking the square root gives (64), from which we have the bound
$\left\|x(t,j)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|\leq M\exp({-\rho}t)\left\|x({0},{0})-\tilde{x}\right\|\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big]\\ +\frac{M\left\|B\right\|q^{\frac{\ell}{2}}d_{\mathcal{U}}}{\rho}\big[1-\exp({-\rho}t)\big].$ (67)
We observe that by definition of $\|\cdot\|_{\mathcal{A}}$ we have
$\|x(0,0)-\tilde{x}\|\leq\|\phi(0,0)\|_{\mathcal{A}}+r.$ (68)
Similarly, from (41) we have
$\|\phi(t,j)\|_{\mathcal{A}}=\|x(t,j)\|_{B_{r}(\tilde{x})}\leq\max\{0,\|x(t,j)-\tilde{x}\|-r\}.$ (69)
Using (68) and (69) in (67) gives
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)(\|\phi(0,0)\|_{\mathcal{A}}+r)\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}[1-\exp({-\rho}t)]\\ +\frac{M\left\|B\right\|q^{\frac{\ell}{2}}d_{\mathcal{U}}}{\rho}[1-\exp({-\rho}t)]-r,$ (70)
where we can remove the $\max$ operator that was in (69) because the right-hand side of (70) is non-negative.
Then we combine like terms to find
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M\exp({-\rho}t)\|\phi(0,0)\|_{\mathcal{A}}\\ +(M-1)\left(\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}+\frac{M\left\|B\right\|q^{\frac{\ell}{2}}d_{\mathcal{U}}}{\rho}\right)\exp({-\rho}t).$ (71)
We find the asymptotic behavior by taking the limit as $t+j$ goes to infinity, which implies that $t$ itself must go to infinity. Then
$\lim_{t+j\to\infty}\exp(-\rho t)=0.$ (72)
Using (72) in (71) gives $\lim_{t+j\to\infty}\|\phi(t,j)\|_{\mathcal{A}}=0$. ∎
Remark 4.
In the limit, the state $x$ is asymptotically no farther than a distance $r$ from $\tilde{x}$. The quantity $r$ can be reduced with a pre-feedback controller that is applied before feedback optimization is used. More precisely, if the pair $(A,B)$ is controllable, then using pole placement we can use any $\eta>0$ to select $\rho$ such that $\rho\geq\frac{M\left\|B\right\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{\mathcal{U}}}}{\eta}(1+q^{\frac{\ell}{2}})$. This $\rho$ ensures $r\leq\eta$ and hence that $\|x(t,j)\|_{B_{\eta}(\tilde{x})}\to 0$ for any desired $\eta>0$.
V-D Global Convergence and Robustness
Theorem 1 relies on (20), which restricts it to only apply to initial conditions with $\tau_{c}(0,0)\in[\tau_{c,min},\tau_{c,max}]$ and $\tau_{g}(0,0)=\tau_{g,comp}$. We next derive a global convergence result that applies to solutions that begin from arbitrary initial conditions, including those that violate the conditions in (20).
Theorem 2 (Global Complete Hybrid Convergence).
Consider the hybrid system $\mathcal{H}_{FO}$ from (17) and suppose that Assumptions 1 and 2 hold. Consider objectives of the form of (8), and suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$, where $Q_{u}$ is from (8) and $L:=\lambda_{max}(Q_{u}+H^{\top}Q_{y}H)$. For each maximal solution $\phi$ to $\mathcal{H}_{FO}$ with initial condition $\phi(0,0)$, for each $(t,j)\in\textnormal{dom }\phi$,
$\left\|\phi(t,j)\right\|_{\mathcal{A}}\leq M^{2}\exp(-\rho\tau_{c,\min})\exp(-\rho t)\left\|\phi(0,0)\right\|_{\mathcal{A}}\\ +\frac{M^{2}\left\|B\right\|d_{\mathcal{U}}}{\rho}\exp(-\rho t)\times\\ \Bigg[\left(M\exp(-\rho\tau_{c,\min})-\frac{1}{M}\right)\Big(1+q^{\frac{\ell}{2}}\Big)+2\Bigg],$ (73)
where $d_{\mathcal{U}}=\max_{u_{1},u_{2}\in\mathcal{U}}\|u_{1}-u_{2}\|$ is the diameter of the set $\mathcal{U}$, $\rho$ is from (38), $\ell\geq 1$ is from Assumption 2, $\mathcal{A}$ is from (36), $q:=1-2\gamma\lambda_{min}(Q_{u})+\gamma^{2}L^{2}\in\left(0,1\right)$, and $M\geq 1$ is a constant.
Proof.
To derive a bound without the conditions in (20), we must consider any $\tau_{c}(0,0)\in[0,\tau_{c,max}]$ and any $\tau_{g}(0,0)\in[0,\tau_{g,comp}]$. The analysis in Theorem 1 does not cover this case because it assumes that there are at least $\ell\geq 1$ gradient descent iterations performed before the first change in the input, but that is no longer guaranteed if the conditions in (20) do not hold. However, after a jump is triggered by $\tau_{c}$ reaching zero, it is guaranteed that at least $\ell$ gradient descent iterations will be performed before the second change in the input. This second change in the input occurs at hybrid time $(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)$, and we focus on bounding the behavior of $\mathcal{H}_{FO}$ until this time.
The input over the interval $t_{0},t_{\bar{\alpha}(1)+1}]$ is $u(0,0)$. To derive a bound that holds from all initial conditions, we consider $\alpha(0)=0$ gradient descent iterations being completed before time $t_{\bar{\alpha}(1)+1}$, which represents the least progress that the underlying optimization algorithm may make before the first change in the input. Then the next input is simply equal to the previous input, and in particular the input over the interval $t_{\bar{\alpha}(1)+1},t_{\bar{\alpha}(2)+2}]$ is
$u(t_{\bar{\alpha}(1)+1},\bar{\alpha}(1)+1)=u(0,0),$ (74)
which happens precisely because no computations have been performed to change the input. This case is captured by using the same steps to reach (67) but with $\ell=0$, which gives
$\left\|x(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|\leq\\ M\exp({-\rho}t_{\bar{\alpha}(2)+2})\left\|x({0},{0})-\tilde{x}\right\|+\frac{2M\left\|B\right\|d_{\mathcal{U}}}{\rho}.$ (75)
To maximize the upper bound, we observe that
$\tau_{c,\min}\leq t_{\bar{\alpha}(2)+2}$ (76)
since, by the point at which flow time $t_{\bar{\alpha}(2)+2}$ is reached, the timer $\tau_{c}$ has undergone at least one reset to the interval $\tau_{c,\min},\tau_{c,\max}]$, followed by $\tau_{c}$ reaching zero. Using (76) in (75) gives
$\left\|x(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{x}}\right\|\leq\\ M\exp({-\rho}\tau_{c,\min})\left\|x({0},{0})-\tilde{x}\right\|+\frac{2M\left\|B\right\|d_{\mathcal{U}}}{\rho}.$ (77)
Then, using (68) and (69) with $(t,j)=(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)$ and combining like terms gives
$\left\|\phi(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)\right\|_{\mathcal{A}}\leq M\exp({-\rho}\tau_{c,\min})\left\|\phi({0},{0})\right\|_{\mathcal{A}}\\ +M\exp({-\rho}\tau_{c,\min})\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(1+q^{\frac{\ell}{2}}\Big)\\ +\frac{M\left\|B\right\|d_{\mathcal{U}}}{\rho}\Big(1-q^{\frac{\ell}{2}}\Big).$ (78)
After hybrid time $(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)$, there are at least $\ell$ gradient descent iterations performed before each subsequent change in the input. Then Theorem 1 applies but with the upper bound on $\|\phi(t_{\bar{\alpha}(2)+2},\bar{\alpha}(2)+2)\|_{\mathcal{A}}$ in place of $\|\phi(0,0)\|_{\mathcal{A}}$. The result follows from making that substitution.
In Theorem 2 one can use the value of $\rho$ in Remark 4 to force $x$ to converge to a ball of any radius $\eta>0$ about $\tilde{x}$.
We can represent modeling errors as perturbations applied to the nominal system $\mathcal{H}_{FO}$. We consider both errors in the LTI dynamics and errors in the timer dynamics. We first consider the perturbed domain of the flow map, which is
$C_{\rho}:=\{\zeta\in\mathcal{X}:\tau_{c}\in[0,\tau_{c,\max}+\theta_{c,\max}],\\ \tau_{g}\in[0,\tau_{g,comp}+\theta_{g,comp}]\},$ (79)
where $\theta_{c,\max}\in(-\tau_{c,\max},\infty)$ and $\theta_{g,comp}\in(-\tau_{g,comp},\infty)$. These perturbations allow $\tau_{c}$ to take values larger than $\tau_{c,\max}$ and similar for $\tau_{g}$. Jumps are still triggered when at least one timer reaches zero and hence we use $D_{\rho}:=D$ in the perturbed system model.
In the flow map, there may be model errors in the $A$ and $B$ matrices, and the two timers may count down at a rate that is not exactly $1$. We define the perturbed flow map as
$F_{\rho}(\zeta):=\left(\begin{array}[]{cc}\big(A+\hat{A}\big)x+\big(B+\hat{B}\big)u\\ 0\\ 0\\ 0\\ -1+\kappa_{c}\\ -1+\kappa_{g}\end{array}\right),$ (80)
where $\hat{A}\in\mathbb{R}^{n\times n}$ models errors in the $A$ matrix, $\hat{B}\in\mathbb{R}^{n\times m}$ models errors in the $B$ matrix, $\kappa_{c}\in(-\infty,1)$ models errors in the rate at which $\tau_{c}$ counts down, and $\kappa_{g}\in(-\infty,1)$ models errors in the rate at which $\tau_{g}$ counts down.
Finally we define the perturbed jump map as
[…]
For $G_{1,\rho}$ we have
$G_{1,\rho}(\zeta):=\left(\begin{array}[]{cc}x\\ u\\ y_{s}\\ \Pi_{\mathcal{U}}\left[z-\gamma\nabla_{u}\Phi(z,y_{s})\right]\\ \tau_{c}\\ \tau_{g,comp}+\theta_{g,comp}\end{array}\right),$ (82)
where $\theta_{g,comp}$ is from (79). This perturbed jump map allows for $\tau_{g}$ to be reset to values other than $\tau_{g,comp}$ after a gradient descent iteration is performed. For $G_{2,\rho}$ we have
$G_{2,\rho}(\zeta):=\left(\begin{array}[]{cc}x\\ z\\ \big(H+\hat{H}\big)u+d\\ z\\ \left[\tau_{c,\min}+\theta_{c,\min},\tau_{c,\max}+\theta_{c,\max}\right]\\ \tau_{g}\end{array}\right).$ (83)
Here, $\hat{H}\in\mathbb{R}^{p\times m}$ models perturbations to the matrix $H$, including those that come from the errors $\hat{A}$ and $\hat{B}$ as described above, as well as errors in the output map $\Psi$. The interval to which $\tau_{c}$ is reset is perturbed with constants $\theta_{c,\min}\in(-\tau_{c,\min},\infty)$ and $\theta_{c,\max}\in(-\tau_{c,\max},\infty)$ that satisfy $0<\tau_{c,\min}+\theta_{c,\min}\leq\tau_{c,\max}+\theta_{c,\max}$, which ensures that $\tau_{c}$ is reset to a non-empty set, though both endpoints of the interval can be perturbed. Finally we have $G_{3,\rho}(\zeta):=G_{1,\rho}(\zeta)\cup G_{2,\rho}(\zeta)$. We define
$\rho(\zeta)=\max\{\theta_{g,comp},\|\hat{A}x\|,\|\hat{B}u\|,\|\hat{H}u\|,\\ \kappa_{c},\kappa_{g},\theta_{c,\min},\theta_{c,\max}\}$ (84)
to be the maximum size of any perturbation at the state $\zeta\in\mathcal{X}:=\mathbb{R}^{n+2m+p+2}$. The perturbed hybrid system model is
$\mathcal{H}_{FO}^{\rho}=\begin{cases}\dot{\zeta}\in F_{\rho}(\zeta)&\zeta\in C_{\rho}\\ \zeta^{+}\in G_{\rho}(\zeta)&\zeta\in D_{\rho}\end{cases}.$ (85)
Our next theorem provides robustness guarantees for $\mathcal{H}_{FO}$. First, we require the following definition.
Definition 2 ($(\tau,\epsilon)$-closeness [18, Definition 5.23]).
Given $\tau,\epsilon>0$, two hybrid arcs $\phi_{1}$ and $\phi_{2}$ are $(\tau,\epsilon)$-close if
  1. 1.
    for all $(t,j)\in\textnormal{dom }\phi_{1}$ with $t+j\leq\tau$ there exists $s$ such that $(s,j)\in\textnormal{dom }\phi_{2}$, $|t-s|<\epsilon$, and
    $|\phi_{1}(t,j)-\phi_{2}(s,j)|<\epsilon;$ (86)
  2. 2.
    for all $(t,j)\in\textnormal{dom }\phi_{2}$ with $t+j\leq\tau$ there exists $s$ such that $(s,j)\in\textnormal{dom }\phi_{1}$, $|t-s|<\epsilon$, and
    $|\phi_{2}(t,j)-\phi_{1}(s,j)|<\epsilon.$ (87)
To the best of our knowledge, the next result is the first analytical characterization of the robustness of feedback optimization in a hybrid or sampled-data setting.
Theorem 3 (Robustness of $\mathcal{H}_{FO}$).
Consider the hybrid system $\mathcal{H}_{FO}^{\rho}$ with $\rho$ as defined in (84), and suppose that Assumptions 1 and 2 hold. Consider objectives of the form of (8), and suppose that the gradient descent algorithm uses a stepsize $\gamma\in\left(0,\frac{2}{\lambda_{min}(Q_{u})+L}\right)$. Then, for every $\epsilon>0$ and $\tau>0$, there exists $\delta>0$ with the following property: for every solution $\phi_{\delta}$ to $\mathcal{H}_{FO}^{\delta\rho}$, there exists a solution $\phi$ to $\mathcal{H}_{FO}$ such that $\phi_{\delta}$ and $\phi$ are $(\tau,\epsilon)$-close.
Proof.
Since $\mathcal{H}_{FO}$ is well-posed and its maximal solutions are complete from all initial conditions, the result follows from [26, Proposition 6.34]. ∎
The interpretation of Theorem 3 is as follows. Let a time $\tau$ and an error $\epsilon$ be given. Then there is a nonzero perturbation $\delta\rho$ that we use to get the perturbed system $\mathcal{H}_{FO}^{\rho\delta}$, and any solution $\phi_{\delta}$ to $\mathcal{H}_{FO}^{\rho\delta}$ is $(\tau,\epsilon)$-close to some trajectory that the unperturbed system $\mathcal{H}_{FO}$ could have produced.
VI Simulation Results
Figure 2: The left-hand plot shows the convergence of the state $x$ under the dynamics $\mathcal{H}_{FO}$ from the initial condition in (90). We see that $x$ attains an asymptotic error of $5\cdot 10^{-9}$, which illustrates that it converges quite close to its desired value. The right-hand plot shows the convergence of the state $x$ under the perturbed dynamics $\mathcal{H}_{FO}^{\rho}$ from the initial condition in (90). In this case, $x$ attains an asymptotic error of $1.52$, even though there are simultaneously perturbations to the timer dynamics and the model of the underlying LTI system.
In this section we present two sets of simulation results: one with the nominal system $\mathcal{H}_{FO}$ and one with the perturbed system $\mathcal{H}_{FO}^{\rho}$. We consider the LTI dynamics
$\dot{x}$ $=\begin{bmatrix}-3&0&0&0\\ 3&-3&0&0\\ 0&3&-3&0\\ 0&0&3&-3\\ \end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\\ x_{4}\end{bmatrix}+\begin{bmatrix}3\\ 0\\ 0\\ 0\\ \end{bmatrix}u$ (88)
$y$ $=x,$
and the feedback optimization problem we solve is
[…]
subject to $y=Hu+d,\,\,u\in\mathcal{U},\,\,y\in\mathbb{R}^{4},$
where we use $Q_{u}=0.08$, $Q_{y}=0.3I_{4}$, $\mathcal{U}=[0,12]$, $d=(0.2,0.2,0.2,0.2)^{\top}$, and $\hat{y}=(4,4,4,4)^{\top}$.
For simulations, the Hybrid Equations Toolbox (Version 3.0.0.76) was used, along with the initial conditions
$x(0,0)=(0,3,10,15)^{\top},u(0,0)=0,\\ y_{s}(0,0)=(0.2,3.2,10.2,15.2)^{\top},z(0,0)=0,\\ \tau_{c}(0,0)=0.175,\penalty 10000\ \tau_{g}(0,0)=0.05,$ (90)
where $\tau_{g,comp}=0.05,\tau_{c,\min}=0.15$, and $\tau_{c,\max}=0.20$ with stepsize $\gamma=0.35\leq\frac{2}{\lambda_{min}(Q_{u})+L}$. We see in the left-hand plot of Figure 2 that $\|x-\tilde{x}\|$ converges to $5\cdot 10^{-9}$, and thus $x$ is asymptotically close to its desired value.
We next consider the perturbed case. Let $J_{a,b}$ denote the matrix of all ones in $\mathbb{R}^{a\times b}$. For the timers, we use perturbations of the form $\theta_{c,\min}=\theta_{c,\max}=\theta_{g,comp}=\kappa_{c}=\kappa_{g}=\chi$ for several values of $\chi$, and for the LTI dynamics we use perturbations of the form $\hat{A}=\mu J_{n\times n}$, $\hat{B}=\mu J_{n\times m}$, and $\hat{H}=\mu J_{p\times m}$ for several values of $\mu$. The values of $\chi$ and $\mu$ used in simulations are shown in Table I, along with the values they induce in the steady-state error $\|x-\tilde{x}\|$. Using these values, the perturbation $\rho$ from (84) can be quite large because it is the maximum of several terms, including $\|\hat{A}x\|$, $\|\hat{B}u\|$, and $\|\hat{H}u\|$, which can be large because they depend on $x$ or $u$.
TABLE I: Values of perturbations and resulting asymptotic error for the perturbed system $\mathcal{H}_{FO}^{\rho}$.
Limiting value of $\|x-\tilde{x}\|$
$\chi=0.2$, $\mu=0.01$ 0.28
$\chi=0.4$, $\mu=0.02$ 0.51
$\chi=0.6$, $\mu=0.03$ 0.21
$\chi=0.8$, $\mu=0.04$ 0.64
$\chi=0.95$, $\mu=0.05$ 0.80
Across all runs, $\|x-\tilde{x}\|$ ranged from $0.28$ to $0.80$, which indicates that $\mathcal{H}_{FO}^{\rho}$ successfully reduces error and rejects disturbances, even under perturbations $\rho$ whose size is comparable to the size of other terms in the dynamics.
VII Conclusion
This paper presented a hybrid system model for feedback optimization that considers continuous-time dynamics with discrete-time optimization. We showed that its maximal solutions are complete and non-Zeno, and then we bounded their distance to a desired goal state. We also presented what are, to the best of our knowledge, the first analytical robustness results for feedback optimization in a hybrid/sampled-data setting. Future work includes using nonconvex objective functions and using hybrid feedback optimization for systems with nonlinear dynamics.